A vector perpendicular to the vector(i^ +2j^) and having magnitude 3√5 units is
 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
let the perpendicular vector = ai + bj
 
now
(ai + bj) . (i + 2j) = 0
a + 2b = 0
 
and 
 
a^2 + b^2 = 45
 
solve these two equations to fid a and b